OpenStudy (anonymous):

the total sales of a company after t weeks is given by the integral: S(t)= 60x^2/sqrt(x+9)dx (it is assumed that the sales are initially zero.) a) find the formula for the sales after t weeks, S(t) b) use your formula to find the total sales afetr 1 year (52 weeks).

OpenStudy (anonymous):

\[\int\limits_{0}^{t}60x^2/\sqrt{x+9}dx\]

OpenStudy (zarkon):

u=x+9

OpenStudy (anonymous):

thanks for the hint